# Learning from Heterogeneity: A Dynamic Learning Framework for Hypergraphs

### v2

@@ V-B Performance Analysis (RQ1)
V-B2 Link Prediction (RQ1.2)
To investigate the effectiveness of our proposed framework, we also apply our model to the link prediction task. In this task, a graph with a certain faction of edges removed is given, and thus the objective is to predict these missing edges. Following the experiment setting adopted in [47], 50% percent of existing edges are randomly set hidden as positive samples while ensuring that the residual graph after the edge being removed remains connected. To generate negative examples, we randomly sample an equal number of node pairs from the graph, which have no edge connecting them. Similar to the related works [16, 48], we use a logistic regression classifier as a predictor for link prediction. To get the representation of each edge, we use Hadamard operator to compute the element-wise product for the linked node pairs. We compare the performance of LFH with five baselines. Node2vec [47] is a classic model for link prediction task. DHGNN is a hypergraph learning method that incorporates dynamic hypergraph construction. The other three baselines, GAT, PC-HGN, and HGNN+, are the top-performing baselines in homogenous pairwise graph learning models, heterogeneous pairwise graph learning models, and hypergraphs learning models in node classification tasks. Experimental results are summarized in Table V. LFH obtains performance gains of 16.8%, 18.8%, 2.9%, 14.6% and 13.7%, compared with Node2vec, GAT, PC-HGN, HGNN+, DHGNN and outperforms five baseline models in three datasets by an average improvement of 12.8%. It shows the promising capability of our model to capture implicit data relations. It is worth noting that HGNN+ that performs well in node classification does not have comparable performance in link prediction. This is because the HGNN+ applied for link prediction task heavily relies on pairwise edges to generate hyperedges, and thus a sharp decrease in the number of existing edges leads to a deterioration in performance.
@@
V-C Impact of Hyperedge Construction (RQ2)
In our proposed LFH, $\lambda$ and $\gamma$ are used as hyperparameters. Specifically, $\lambda$ is the weight hyperparameter of the reconstruction error and $\gamma$ is the norm hyperparameter to trade off $l_{1}\text{-}norm$ regularization norm and $l_{2}\text{-}norm$ regularization of the reconstruction coefficient vector (See Eq. 6). Figure 4 and 5 illustrate the overall performance of LFH influenced by $\lambda$ and $\gamma$. Figure 4 illustrates the influence of $\lambda$ for different datasets on F1 score. In this figure, we fix $\gamma$ as 0.02, 0.2 and 0.2 while $\lambda$ ranges between 0.0001 and 1. When $\lambda$ is small, the reconstruction loss of $\lambda c^{k}\left(v^{m}_{i}\right)$ becomes trivial, and the reconstruction coefficient vector p will not experience large adjustments throughout the training phase, thus affecting the generation of hyperedges and deteriorating the performance of LFH. In contrast, when $\lambda$ increases to around 0.2, reconstruction loss of $\lambda c^{k}\left(v^{m}_{i}\right)$ accounts for the appropriate proportion of the total loss, thus achieving the best performance under this setup. When $\lambda$ continually increases to 20, F1 score drops rapidly, which clearly demonstrates that the dominance of reconstruction loss compromises the generalization ability of LFH.
@@ V-C Impact of Hyperedge Construction (RQ2)
V-D Sensitivity Analysis (RQ3)
We study the impact of key parameters in LFH including the node embedding size and the number of attention head as illustrated in Table VI and VII.
V-D1 Impact of the node embedding size
We test the impact of different node embedding sizes for all datasets with data split ratio of 10%. As shown in Table VI, the performance of the model increases at first and then starts to drop, reporting the best result with the embedding size 256. Intuitively, larger embedding sizes could lead to overfitting, thus causing unexpected performance drops.
V-D2 Impact of the number of muti-head
We analyse the effect of the number of multi-head $K$. As observed in Table VII, when the number of attention heads increases, there is typically an improvement in the performance of our model. However, it is observed that the performance of LFH improves only marginally afterwards while causing large computational costs. When $K=4$, it is an optimal point to trade off the performance and computational cost.
TABLE VI: Performance of LFH with different node embedding sizes.
@@ V-D2 Impact of the number of muti-head
TABLE VII: Performance of LFH with different number of attention heads.
@@
Figure 6: Ablation on using different GNN models for initial embedding generation
TABLE VIII: Performance of LFH for ablation study.
@@
V-E Ablation Study (RQ4)
We study the impact of pairwise fusion in this part. We consider different candidates in the pairwise fusion process, aiming to find the best match that is capable of deriving high-quality features from the heterogeneous pairwise graph as the initial node embedding. We analyse the importance of pairwise fusion and implement different models introduced in Section V-A. Note that we applied the same hyperparameters of all these models according to the best performance reported in the respective papers and source codes. As shown in Figure 6, using PC-HGN in pairwise fusion enlarges the performance gap over other models in all datasets, with a performance gain of more than 5.4%-14.4% compared to using GAT, a performance gain of more than 6.6%-18.0% compared to using HGT, and an overall 11.5% performance gain on average compared to using other models in pairwise fusion. This result demonstrates the initial node embedding generated by PC-HGN helps our framework achieve the best results in all three datasets.
In order to validate the effectiveness of different components in LFH, we design two variants of LFH, and compare the performance of these variants with LFH in the node classification task. The results of comparison are shown in Table VIII. Specifically, we investigate the contributions of dynamic hypergraph construction and multi-head attention to the performance of LFH. LFH w/o dynamic hypergraph construction replaces the learning process with unsupervised $k$-NN method for generating heterogeneous hyperedges, and the number of nearest nodes to form a hyperedge is set to 10 [19]. The results show that LFH has a better performance over LFH w/o dynamic hypergraph construction by margins of 4.4%-18.9%, which indicates that the learning process for generating hypergraph enables LFH to encode topological structure more accurately. LFH w/o multi-head attention excludes the importance of heterogeneous hyperedges with respect to the master node, and assigns the weights to different heterogeneous hyperedges equally during hypergraph learning. The performance of LFH w/o multi-head attention drops significantly in all three datasets. The result demonstrates the significance of the dynamically learned weights of hyperedges and the effectiveness of the multi-head attention mechanism.
@@ V-E Ablation Study (RQ4)
V-F Impact of Hyperparameter $\alpha$ (RQ5)
@@
V-G Visualization
@@
To investigate the effectiveness of our model compared to other graph learning models in encoding the heterogeneity of the topology, we visualize the learned node embeddings using the t-SNE tool [49] on the ACM dataset. It is worth noting that ACM dataset naturally possess heterogeneous properties, as represented by various types of nodes (such as Author, Paper and Subject in ACM). Visualizing the distributions of nodes with different types directly demonstrates the method’s ability to capture heterogeneity. We compare LFH with other three baselines, GAT, PC-HGN, and HGNN+. These models represent the top-performing baselines for node classification tasks in homogeneous pairwise graph learning, heterogeneous pairwise graph learning, and hypergraph learning. As shown in Fig. 8, our proposed model LFH yields discernible clusters compared to other baseline models on the ACM dataset, which verifies the effectiveness of the proposed method in capturing the heterogeneity of the topology.
VI conclusion
@@
In this paper, we propose a heterogenous hypergraph learning framework for representation learning. This framework first generates the high-quality initial node embedding using the designated pairwise fusion function, aiming to exploit the pairwise graph information at most. Afterwards, the multi-type hyperedges are constructed dynamically, forming the hypergraph together. The embedding is then updated iteratively through type-specific attention, aiming to encode the heterogeneous attributes into the embedding space. We conduct comprehensive experiments on three widely-used public datasets and comparison with twelve baseline methods to demonstrate the effectiveness of our proposed framework. The results and analysis reveal that the proposed framework can achieve new state-of-the-art performance on both node classification and link prediction tasks.

### v3

@@ 5.2 Performance Analysis (RQ1)
5.2.2 Link Prediction (RQ1.2)
To investigate the effectiveness of our proposed framework, we also apply our model to the link prediction task. In this task, a graph with a certain faction of edges removed is given, and thus the objective is to predict these missing edges. Following the experiment setting adopted in [48], 50% percent of existing edges are randomly set hidden as positive samples while ensuring that the residual graph after the edge being removed remains connected. To generate negative examples, we randomly sample an equal number of node pairs from the graph, which have no edge connecting them. Similar to the related works [16, 49], we use a logistic regression classifier as a predictor for link prediction. To get the representation of each edge, we use Hadamard operator to compute the element-wise product for the linked node pairs. We compare the performance of LFH with five baselines. Node2vec [48] is a classic model for link prediction task. DHGNN is a hypergraph learning method that incorporates dynamic hypergraph construction. The other three baselines, GAT, PC-HGN, and HGNN+, are the top-performing baselines in homogenous pairwise graph learning models, heterogeneous pairwise graph learning models, and hypergraphs learning models in node classification tasks. Experimental results are summarized in Table 5. LFH obtains performance gains of 16.8%, 18.8%, 2.9%, 14.6% and 13.7%, compared with Node2vec, GAT, PC-HGN, HGNN+, DHGNN and outperforms five baseline models in three datasets by an average improvement of 12.8%. It shows the promising capability of our model to capture implicit data relations. It is worth noting that HGNN+ that performs well in node classification does not have comparable performance in link prediction. This is because the HGNN+ applied for link prediction task heavily relies on pairwise edges to generate hyperedges, and thus a sharp decrease in the number of existing edges leads to a deterioration in performance.
@@
5.3 Impact of Hyperedge Construction (RQ2)
In our proposed LFH, $\lambda$ and $\gamma$ are used as hyperparameters. Specifically, $\lambda$ is the weight hyperparameter of the reconstruction error and $\gamma$ is the norm hyperparameter to trade off $l_{1}\text{-}norm$ regularization norm and $l_{2}\text{-}norm$ regularization of the reconstruction coefficient vector (See Eq. 6). Figure 5 and 6 illustrate the overall performance of LFH influenced by $\lambda$ and $\gamma$. Figure 5 illustrates the influence of $\lambda$ for different datasets on F1 score. In this figure, we fix $\gamma$ as 0.02, 0.2 and 0.2 while $\lambda$ ranges between 0.0001 and 1. When $\lambda$ is small, the reconstruction loss of $\lambda c^{k}\left(v^{m}_{i}\right)$ becomes trivial, and the reconstruction coefficient vector p will not experience large adjustments throughout the training phase, thus affecting the generation of hyperedges and deteriorating the performance of LFH. In contrast, when $\lambda$ increases to around 0.2, reconstruction loss of $\lambda c^{k}\left(v^{m}_{i}\right)$ accounts for the appropriate proportion of the total loss, thus achieving the best performance under this setup. When $\lambda$ continually increases to 20, F1 score drops rapidly, which clearly demonstrates that the dominance of reconstruction loss compromises the generalization ability of LFH.
@@ 5.3 Impact of Hyperedge Construction (RQ2)
5.4 Sensitivity Analysis (RQ3)
We study the impact of key parameters in LFH including the node embedding size and the number of attention head as illustrated in Table 6 and 7.
5.4.1 Impact of the node embedding size
We test the impact of different node embedding sizes for all datasets with data split ratio of 10%. As shown in Table 6, the performance of the model increases at first and then starts to drop, reporting the best result with the embedding size 256. Intuitively, larger embedding sizes could lead to overfitting, thus causing unexpected performance drops.
5.4.2 Impact of the number of muti-head
We analyse the effect of the number of multi-head $H$. As observed in Table 7, when the number of attention heads increases, there is typically an improvement in the performance of our model. However, it is observed that the performance of LFH improves only marginally afterwards while causing large computational costs. When $H=4$, it is an optimal point to trade off the performance and computational cost.
Table 6: Performance of LFH with different node embedding sizes.
@@ 5.4.2 Impact of the number of muti-head
Table 7: Performance of LFH with different number of attention heads.
@@
Figure 7: Ablation on using different GNN models and random initialization strategies for initial embedding generation.
Table 8: Performance of LFH for ablation study.
@@
5.5 Ablation Study (RQ4)
We study the impact of pairwise fusion in this part. We consider different candidates in the pairwise fusion process, aiming to find the best match that is capable of deriving high-quality features from the heterogeneous pairwise graph as the initial node embedding. We analyse the importance of pairwise fusion and implement different models introduced in Section 5.1. Note that we applied the same hyperparameters of all these models according to the best performance reported in the respective papers and source codes. As shown in Figure 7, using PC-HGN in pairwise fusion to generate initial node embeddings significantly enhances performance across all datasets. It yields a performance gain of more than 5.4% to 14.4% compared to GAT, more than 6.6% to 18.0% compared to HGT, and an average performance gain of 10.5% compared to the random initialization strategy. This result further verify the great importance to generate high-quality initial node embedding to better exploit the implicit data relations in the hypergraph. Additionally, it demonstrates the initial node embedding generated by PC-HGN helps our framework achieve the best results in all three datasets.
In order to validate the effectiveness of different components in LFH, we design two variants of LFH, and compare the performance of these variants with LFH in the node classification task. The results of comparison are shown in Table 8. Specifically, we investigate the contributions of dynamic hypergraph construction and multi-head attention to the performance of LFH. LFH w/o dynamic hypergraph construction replaces the learning process with unsupervised $k$-NN method for generating heterogeneous hyperedges, and the number of nearest nodes to form a hyperedge is set to 10 [19]. The results show that LFH has a better performance over LFH w/o dynamic hypergraph construction by margins of 4.4%-18.9%, which indicates that the learning process for generating hypergraph enables LFH to encode topological structure more accurately. LFH w/o multi-head attention excludes the importance of heterogeneous hyperedges with respect to the master node, and assigns the weights to different heterogeneous hyperedges equally during hypergraph learning. The performance of LFH w/o multi-head attention drops significantly in all three datasets. The result demonstrates the significance of the dynamically learned weights of hyperedges and the effectiveness of the multi-head attention mechanism.
@@ 5.5 Ablation Study (RQ4)
5.6 Impact of Hyperparameter $\alpha$ (RQ5)
@@
Figure 10: A t-SNE visualization of our proposed model compared to baseline models on the IMDB dataset (red indicates actor nodes, green denotes movie nodes, and blue represents director nodes).
Figure 11: A t-SNE visualization of our proposed model compared to baseline models on the DBLP dataset (red indicates paper nodes, green denotes author nodes, and blue represents conference nodes).
5.7 Visualization
@@
To investigate the effectiveness of our model compared to other graph learning models in encoding the heterogeneity of the topology, we visualize the learned node embeddings using the t-SNE tool [50] on the ACM, IMDB and DBLP datasets. It is worth noting that these three datasets naturally possess heterogeneous properties, as represented by various types of nodes (such as Author, Paper and Subject in ACM). Visualizing the distributions of nodes with different types directly demonstrates the method’s ability to capture heterogeneity. We compare LFH with other three baselines, GAT, PC-HGN, and HGNN+. These models represent the top-performing baselines for node classification tasks in homogeneous pairwise graph learning, heterogeneous pairwise graph learning, and hypergraph learning. As shown in Fig. 9-11, our proposed model LFH yields discernible clusters compared to other baseline models across the ACM, IMDB and DBLP datasets, which verifies the effectiveness of the proposed method in capturing the heterogeneity of the topology.
Figure 12: (a) Runtime cost with different baseline models on three datasets. (b) Model performance comparison.
5.8 Timing Efficiency Analysis
We compare the time efficiency on a workstation equipped with an Intel Core i7 CPU @ 2.6 GHz and 32 GB of RAM against four baseline models. DHGNN is a hypergraph model utilizing the cluster method to dynamically construct the hypergraph. The other three baselines, GAT, PC-HGN, and HGNN+, are the top-performing baselines in homogenous pairwise graph learning models, heterogeneous pairwise graph learning models, and hypergraphs learning models in node classification tasks. We set the number of training epochs for all models to 100 to make a fair comparison across three datasets, and we report the results for runtime cost and performance in Fig. 12.
From Fig. 12 (a), it is observed that LFH achieves a significantly lower training time cost compared to DHGNN and PC-HGN on all three datasets. It is worth noting that the runtime reported for our model excludes the time spent on initial node embedding. The encoded pairwise connectivities in pre-trained initial node embedding provide our model with a better starting point from the beginning of the training and thus enhance the learning of representation. As illustrated in Fig. 12 (b), LFH demonstrates superior performance on all three datasets when compared to all competing baseline models. Apart from that, we observe that both HGNN+ and GAT require less time for training than our model. However, despite time efficiency in the training process, HGNN+ and GAT both fail to capture the heterogeneity, leading to an inferior performance when compared with LFH. The capacity of different models to encode the heterogeneity has been further analyzed in the Section 5.7.
6 conclusion
@@ 5.8 Timing Efficiency Analysis
In this paper, we propose a heterogenous hypergraph learning framework for representation learning. This framework first generates the high-quality initial node embedding using the designated pairwise fusion function, aiming to exploit the pairwise graph information at most. Afterwards, the multi-type hyperedges are constructed dynamically, forming the hypergraph together. The embedding is then updated iteratively through type-specific attention, aiming to encode the heterogeneous attributes into the embedding space. We conduct comprehensive experiments on three widely-used public datasets and comparison with twelve baseline methods to demonstrate the effectiveness of our proposed framework. The results and analysis reveal that the proposed framework can achieve new state-of-the-art performance on both node classification and link prediction tasks. In the future, we plan to extend our method to tackle regression problems involving temporal-spatial data by constructing hyperedges that incorporate temporal information, thereby better encoding the evolution of embeddings from this data. Furthermore, we also intend to apply our method to other fields where data naturally have a graph structure, such as protein data in bioinformatics, to provide better representative embeddings for downstream tasks.